# Giant and robust Josephson diode effect in multiband topological nanowires

### v2

@@ Abstract
Introduction.– The Majorana bound states (MBSs) have attracted significant interest for obeying non-Abelian statistics, and can serve as a basic building block for fault tolerant quantum computing Alicea (2012); Kitaev (2003); Nayak et al. (2008); Ivanov (2001); He et al. (2020); Chan et al. (2017). A leading platform for realizing MBSs is the proximitized nanowire Lutchyn et al. (2010); Oreg et al. (2010); Sau et al. (2010); Laubscher and Klinovaja (2021); Stanescu et al. (2011); Lutchyn et al. (2011); Qiao et al. (2024), where a semiconductor wire with strong spin-orbit interaction (SOI) is coupled to a superconductor under a magnetic field. This setup can host MBSs at two ends of the nanowire, with experimental signatures such as zero-bias conductance peaks Prada et al. (2020); Sasaki et al. (2011); Deng et al. (2012); Das et al. (2012); Mourik et al. (2012); Churchill et al. (2013) and the fractional $4\pi$-periodic Josephson effect Kwon et al. (2004a); Fu and Kane (2009); Kwon et al. (2004b); Laroche et al. (2019); Wiedenmann et al. (2016). The latter provides a key tool for probing MBSs, reflecting their unique phase behavior in Josephson junctions. Despite significant progress Wang et al. (2018); Albrecht et al. (2016); Deng et al. (2016); Quantum et al. (2025); Ren et al. (2019), obtaining unambiguous confirmation of MBSs remains a persistent challenge that continues to drive ongoing research.
@@
Diodes, which allow current to flow mainly in one direction, are essential in modern electronics and have played a key role in the development of the information age. While conventional diodes suffer from Joule heating due to resistance, superconducting diodes provide a dissipation free alternative. This advantage has spurred rapid interest in the superconducting diode effect (SDE). The SDE is generally divided into intrinsic Ando et al. (2020); Miyasaka et al. (2021); Yuan and Fu (2021); He et al. (2022); Narita et al. (2022); Yuan and Fu (2022a); Daido et al. (2022); Zinkl et al. (2022); Hou et al. (2023); Anh et al. (2024); Mao et al. (2024a) and Josephson diode effects (JDE) Hu et al. (2007a); Misaki and Nagaosa (2021a); Wang et al. (2022); Davydova et al. (2022a); Wu et al. (2022a); Baumgartner et al. (2022a); Pal et al. (2022a); Zhang et al. (2022), with the latter drawing particular attention. When both inversion and time-reversal symmetries are broken Liu and Lobos (2013); Liu (2012), the forward ($I_{c}^{+}$) and reverse ($I_{c}^{-}$) directional critical currents are generally unequal Liu and Lobos (2013), a key mechanism underlying the SDE. In JDE, the Andreev bound states (ABSs) in the junction region contribute to the supercurrent Sauls (2018); Mizushima and Machida (2018); Hays et al. (2018); Tosi et al. (2019); Nichele et al. (2020). Recent experiments have observed the JDE in various systems Ciaccia et al. (2023); Banerjee et al. (2023); Matsuo et al. (2023); Trahms et al. (2023); Pal et al. (2022b); Baumgartner et al. (2022b); Díez-Mérida et al. (2023); Wu et al. (2022b), sparking significant interest. Alongside these experimental advancements, theoretical frameworks have been developed for the JDE behavior Cheng and Sun (2023); Davydova et al. (2022b); Hu et al. (2007b); Misaki and Nagaosa (2021b); Lu et al. (2023); Souto et al. (2022); Halterman et al. (2022); Kokkeler et al. (2022); Fominov and Mikhailov (2022); Kopasov et al. (2021); Gupta et al. (2023); Sun et al. (2023); Jiang et al. (2022); Tanaka et al. (2022); Maiani et al. (2023); Hu et al. (2023); Misaki and Nagaosa (2021b); Steiner et al. (2023); Souto et al. (2022); Pillet et al. (2023); Hess et al. (2023); Costa et al. (2025); Fukaya et al. (2024); Debnath and Dutta (2025), of particular interest is the JDE realized in topological nanowires hosting MBSs. Previous studies have focused on ideal one-dimensional (1D) models Liu et al. (2024); Legg et al. (2023); Mondal et al. (2025), whereas realistic systems with multiband Stanescu et al. (2011); Lutchyn et al. (2011); Liu et al. (2012) have been largely overlooked. Such systems may exhibit distinct mechanisms that can significantly modify the diode efficiency.
[…]
@@
The multiband model.–We start with the multiband model for the quasi-1D nanowire placed on the $s$-wave superconductor to generate the proximity-induced superconductivity, forming a Josephson junction as depicted in Fig. 1(a). The system has finite dimensions $L_{x}$, $L_{y}$, and $L_{z}$ along the three spatial directions, which can be effectively modeled using square or harmonic trapping potentials. In typical nanowire material, the system is strongly confined along the $y$ and $z$ directions. To facilitate the theoretical description, we consider the rectangular cross section with a width of $L_{y}\sim 10^{2}\,\text{nm}$ and a thickness of $L_{z}\sim 10^{0}\,\text{nm}$), resulting in a reasonable approximation of $L_{z}\ll L_{y}\ll L_{x}$. Consequently, we limit our analysis to the lowest eigenstates in the $z$-direction, setting $N_{z}=1$ Stanescu et al. (2011); Lutchyn et al. (2011). As for the $y$-direction, we focus on cases that a few transverse eigenmodes in this direction can be occupied (e.g. with the number of orbits $N_{y}\leq 3$, see Fig. 1(b)). The resulting effective lattice Hamiltonian takes the form $H_{\mathrm{nw}}=H_{0}+H_{\text{SOI}}+H_{B}+H_{\Delta}$, where $H_{\mathrm{SOI}}$ denotes the Rashba SOI, $H_{B}$ accounts for the Zeeman splitting due to an external magnetic field, and $H_{\Delta}$ represents the induced pairing term. The explicit form of each contribution is given by
@@
with electron creation operator denoted by the spinor $c_{\boldsymbol{i}}^{\dagger}=(c_{\boldsymbol{i}\uparrow}^{\dagger},c_{\boldsymbol{i}\downarrow}^{\dagger})$ and $\boldsymbol{i}=(i_{x},i_{y})$. Note that the index $i_{x}$ denotes the real site position along the $x$-direction, while $i_{y}=1,2,...,N_{y}$ is an effective position denoting transverse modes in the $y$ direction and characterizes an orbital degree of freedom. The Hamiltonian consists of four contributions: The free-particle term $H_{0}$ describes nearest-neighbor hopping $t_{x}$ along $\boldsymbol{\delta}_{x}=(1,0)$ and includes an $i_{y}$-site-dependent (actually transverse-mode dependent) chemical potential $\mu_{\boldsymbol{i}}=\mu_{0}-i_{y}^{2}(\pi\hbar)^{2}/(2m^{*}L_{y}^{2})$ Further details are provided in the Supplementary Material , where the energy of transverse modes is approximated by that in a 1D infinite square well Lutchyn et al. (2011), and $m^{*}$ is the effective mass of the electron. The SOI term $H_{\text{SOI}}$ incorporates the longtitutional SOI with intensity $\alpha_{x}$ and the effective transverse SOI between different orbits with intensity $\alpha_{y}\propto L_{y}^{-1}$ , which is nonzero only when $\boldsymbol{\delta}_{y}$=(0,1) Further details are provided in the Supplementary Material , with the Pauli matrices $\hat{\sigma}_{i}(i=x,y,z)$ on spin space. The term $H_{B}$ describes a uniform magnetic field $\boldsymbol{h}=(h_{x},h_{y})$ that applied throughout the entire nanowire, where $h_{x}$ opens a Zeeman gap in the spectrum, while $h_{y}$ shifts the Fermi surface along the $x$ direction, breaking inversion symmetry Liu and Lobos (2013); Yuan and Fu (2022b). Finally, $s$-wave pairing is depicted by $H_{\Delta}$, with $\Delta_{\boldsymbol{i}}$ nonzero only in the superconducting leads, carrying a phase difference $\phi$ across the junction: $\Delta_{\mathrm{R}}=\Delta_{\mathrm{L}}e^{i\phi}=\Delta e^{i\phi}$.
Balance between fractional and conventional Josephson currents.– In the multiband regime, the Fermi level crosses multiple Fermi points, such that the topological nanowire may generically host both MBSs and ABSs in the boundary [Fig. 1(a)], which in the Josephson junction give rise fractional ($4\pi$-periodic) and conventional ($2\pi$-periodic) Josephson current components, denoted $I_{4\pi}$ and $I_{2\pi}$ [Fig. 1(d)], respectively. The total supercurrent at zero temperature versus superconducting phase difference $\phi$ reads Beenakker (1991)
@@
We refer to realistic parameters from experiments on InAs and InSb nanowires placed on Nb or Al superconductors Lutchyn et al. (2011). The nanowire width is $L_{y}=130\,\text{nm}$ and the effective lattice constant in the $x$-direction $a_{x}$ is taken to be $4\,\text{nm}$. Then $m^{*}=0.04m_{e}$, with SOI strength $\alpha_{x}=0.1\,\text{eV}\cdot\mathring$$\text${A}} and BCS pairing potential $\Delta=2.4\,\text{K}$ ($0.2\,\text{meV}$). The corresponding tight-binding parameters for numerical study are $t_{x}=200\Delta$, $\alpha_{x}=6\Delta$, $\alpha_{y}=1.3\Delta$, and let $\Delta=1$ for convenience. Further, to enable a deep topological phase regime for our study, the Fermi surface can be adjusted to be halfway between the two eigenmodes at $k_{x}=0$, which determines the value of $\mu_{0}$. We consider a short junction limit with a normal region length $N_{\text{N}}=2$, and the Josephson current is computed using the recursive Green’s function method Song et al. (2016); Mao et al. (2024b):
@@
Fig. 2 shows the superconducting diode efficiency versus the magnetic field $h_{x}$ for different system parameters like the SOI strength $\alpha_{x}$ (Fig. 2(a)), the nanowire’s transverse length $L_{y}$ (Fig. 2(b)), and $h_{y}$ (Fig. 2(c)). For the single band regime (see the colored dashed lines), we can find that the diode efficiency $\eta$ consistently exhibits a pronounced peak near the critical topological phase transition boundary $h_{x}^{c}=\sqrt{\Delta^{2}-h_{y}^{2}}$ Liu et al. (2024) and diminishes if tuning into deep topological and trivial regimes. In the trivial (topological) regime, the JDE are primarily contributed from the higher harmonic terms of conventional (fractional) Josephson currents $I_{2\pi}$ ($I_{4\pi}$) due to the ABSs (MBSs) Legg et al. (2023), for which the diode effect is weak. Only near the phase transition boundary, which enables a competition between the two types of Josephson currents, an enhanced Josephson diode effect is resulted.
The multiband regime shows a novel new feature that the high diode efficiency is generically obtained in the deep topological phase regime. Within the multiband topological phase, the contributions from ABSs and MBSs coexist, leading to a persistent competition between $I_{4\pi}$ and $I_{2\pi}$ across the entire topological regime. Consequently, the regime of high diode efficiency extends over most of the topological phase region [Fig. 2(a)-(c)]. Moreover, the maximum position of diode efficiency varies depending on system parameters, but is always obtained from the existing balance between $I_{4\pi}$ and $I_{2\pi}$ in deep topological phase. In Fig. 2(d), increasing the chemical potential drives the Fermi level across one, two, and three subbands, defining regimes I, II, and III. Region $\mathrm{I}$ features a low efficiency platform contributed from MBSs and region $\mathrm{II}$ exhibits weak diode effect dominated by ABSs. In contrast, in region $\mathrm{III}$ a high diode efficiency is driven by the coexistence of MBSs and ABSs.
@@
Figure 3: High diode efficiency plateau by spin-parity subband exchanges mechanism. (a) Energy bands before, during, and after the spin-parity band exchange in the 2-orbit ($N_{y}=2$) model with 3 occupied subbands. Colored bands denote different spin-parity subbands, and colored arrows mark the Fermi point shifts induced by the inversion-symmetry-breaking field $h_{y}$. (b) Diode efficiency $\eta$ as a function of $h_{x}$, where the background colors correspond to the band structures shown in (a). (c) Energy bands showing the spin-parity band exchange in the three-orbital ($N_{y}=3$) model with five occupied subbands. (d) $\eta$ versus $h_{x}$, with color segments matching the band structures in (c). Parameters: $\alpha_{x}=10,\,h_{y}=0.8$.
@@
Spin-parity band exchange mechanism.–A more intriguing discovery is that we predict a robust high diode efficiency plateau with a novel spin-parity band exchange mechanism that is obtained only in the multiband regime. The spin-parity $\mathbb{P}_{s}$ of a certain subband is defined as the sign of the eigenvalue of the spin-coupled term of the Hamiltonian $h_{x}\hat{\sigma}_{x}+(h_{y}+2\alpha_{x}\sin k_{x}a_{x})\hat{\sigma}_{y}$ at $k_{y}=0$, where the interband SOI vanishes. As illustrated in Fig. 3, the bands with $\mathbb{P}_{s}=+1$ and $\mathbb{P}_{s}=-1$ are plotted in red and blue color, respectively. The relative positions of normal subbands (with $\Delta=0$) with different spin parities can be tuned by $h_{x}$ and shows a crucial impact on the diode effect. As $h_{x}$ increases, the subbands with opposite spin parities shift oppositely in energy, leading to the spin parity exchange of the two types of subbands when $h_{x}$ exceeds a threshold. We consider first the topological phase for $N_{y}=2$, with three normal subbands crossing Fermi energy [Fig. 3(a)]. We can see that at $h_{x}=0.7\Delta$ with $h_{y}=0.8\Delta$, on Fermi level there are two blue bands with $\mathbb{P}_{s}=-1$ and one red band with $\mathbb{P}_{s}=+1$. At $h_{x}\approx|\mu_{1}-\mu_{2}|/2\approx 3.5\Delta$, the upper two subbands begin to swap positions, and the full exchange is completed at $h_{x}\geq 4.7\Delta$. Then the spin parity configuration is robust for further increasing $h_{x}$. The diode efficiency in Fig. 3(b) highlights three distinct regions corresponding to the band exchange shown on the left. Prior to the band exchange, the diode efficiency remains to be low. At the band exchange region, a sharp enhancement in the diode efficiency emerges. After the complete exchange of the two subbands with different spin parities, a robust high diode efficiency plateau is obtained.
@@
The predicted high diode efficiency plateau is not limited to $N_{y}=2$, but a generic result applicable to the multiband regime. To illustrate this, we consider further the case of $N_{y}=3$, in which the topological phase is obtained by setting the Fermi level to cross five subbands, including two red subbands and three blue subbands [Fig. 3(c)]. With increasing $h_{x}$, two spin parity exchange transitions are obtained, respectively at $h_{x}\approx 6.2\Delta$ and $h_{x}\approx 8.8\Delta$. Around the transition points, the diode efficiency exhibits local maximum or minimum values depending on details. In comparison, after the second transition, the spin parity configuration is stabilized, and the robust high diode efficiency plateau is generically obtained [Fig. 3(d)].
@@
The giant diode efficiency plateau implies a novel mechanism featured by the spin parity configuration of the subbands. As illustrated in Fig. 3(a), the Fermi level intersects three subbands. For $h_{y}=0$, the bands remain symmetric about $k_{x}=0$. With $h_{y}\neq 0$, the 1D dispersions take the form $E_{i_{y}}^{\pm}(k_{x})=2t_{x}(1-\cos k_{x}a_{x})-\mu_{i_{y}}\pm\sqrt{h_{x}^{2}+(h_{y}+2\alpha_{x}\sin k_{x}a_{x})^{2}},$ with the orbit and spin-parity being labeled by $i_{y}$ and $\pm$, respectively. Therefore, upon applying $h_{y}$, the net shift of the two Fermi points in a given subband is
@@
This expression quantifies the directional deformation of Fermi points, where the red bands ($\mathbb{P}_{s}=+1$) shift leftward and the blue bands ($\mathbb{P}_{s}=-1$) shift rightward [Fig. 3(a)]. Fig. 4(a,b) shows the Fermi point positions and their shifts for the three occupied bands with $N_{y}=2$ as a function of the magnetic field $h_{x}$. We find that $(\Delta k_{\textrm{F}})_{1}^{+}$ and $(\Delta k_{\textrm{F}})_{2}^{-}$ initially decrease, cross zero, and eventually saturate. The zero crossing signifies an exchange of the Fermi points. These Fermi point shifts are induced by the transverse field $h_{y}$. A key feature is that the dimensionless ratio $\xi=2|(\Delta k_{F})_{1}^{+}|/[(\Delta k_{F})_{1}^{-}+(\Delta k_{F})_{2}^{-}]$, which quantifies the relative shift between left- and right-shifted total Fermi momenta, saturates to a stable plateau after the spin-parity exchange [the yellow line in Fig. 4(b)]. This saturation corresponds to the emergence of high diode efficiency plateau. In this regime, the two blue bands with spin-parity $\mathbb{P}_{s}=-1$ exhibit nearly identical and constant Fermi point displacements, which mainly contribute the conventional Josephson current $I_{2\pi}$. In contrast, the remaining red subband predominantly contributes to the fractional Josephson current $I_{4\pi}$ related to MBSs. The saturation of $\xi$ signals the balance between the two types of Josephson currents and the high diode efficiency plateau.

### v3

@@ Abstract
Introduction.– The Majorana bound states (MBSs) have attracted significant interest for obeying non-Abelian statistics, and can serve as a cornerstone for fault tolerant quantum computing Alicea (2012); Kitaev (2003); Nayak et al. (2008); Ivanov (2001); He et al. (2020); Chan et al. (2017). A leading platform for realizing MBSs is the proximitized nanowire Lutchyn et al. (2010); Oreg et al. (2010); Sau et al. (2010); Laubscher and Klinovaja (2021); Stanescu et al. (2011); Lutchyn et al. (2011); Qiao et al. (2024), where a semiconductor wire with strong spin-orbit interaction (SOI) is coupled to a superconductor (SC) under magnetic field. This setup can host MBSs at two ends of the nanowire, with experimental signatures such as zero-bias conductance peaks Prada et al. (2020); Sasaki et al. (2011); Deng et al. (2012); Das et al. (2012); Mourik et al. (2012); Churchill et al. (2013) and the fractional $4\pi$-periodic Josephson effect Kwon et al. (2004b); Fu and Kane (2009); Kwon et al. (2004a); Laroche et al. (2019); Wiedenmann et al. (2016). The latter provides a key tool for probing MBSs, reflecting their unique phase behavior in Josephson junctions. Despite significant progress Wang et al. (2018); Albrecht et al. (2016); Deng et al. (2016); Quantum et al. (2025); Ren et al. (2019), obtaining unambiguous confirmation of MBSs remains a persistent challenge that continues to drive ongoing research.
@@
Diodes, which allow current to flow mainly in one direction, are essential in modern electronics. While conventional diodes suffer from Joule heating due to resistance, superconducting diodes provide a dissipation free alternative. This advantage has spurred rapid interest in the superconducting diode effect (SDE). The SDE is generally divided into intrinsic Bauriedl et al. (2022); Ando et al. (2020); Miyasaka et al. (2021); Yuan and Fu (2021); He et al. (2022); Narita et al. (2022); Yuan and Fu (2022b); Daido et al. (2022); Zinkl et al. (2022); Hou et al. (2023); Anh et al. (2024); Mao et al. (2024a); Le et al. (2024); Wan et al. (2024) and Josephson diode effects (JDE) Hu et al. (2007b); Misaki and Nagaosa (2021a); Wang et al. (2022); Davydova et al. (2022b); Wu et al. (2022a); Baumgartner et al. (2022b); Pal et al. (2022a); Guo et al. (2025); Zhang et al. (2022), with the latter drawing particular attention. When both inversion and time-reversal symmetries are broken Liu and Lobos (2013); Liu (2012), the forward ($I_{c}^{+}$) and reverse ($I_{c}^{-}$) directional critical currents are generally unequal Liu and Lobos (2013), a key mechanism underlying the SDE. In JDE, the Andreev bound states (ABSs) in the junction region contribute to the supercurrent Sauls (2018); Mizushima and Machida (2018); Hays et al. (2018); Tosi et al. (2019); Nichele et al. (2020). Recent experiments have observed the JDE in various systems Ciaccia et al. (2023); Banerjee et al. (2023); Matsuo et al. (2023); Trahms et al. (2023); Pal et al. (2022b); Baumgartner et al. (2022a); Díez-Mérida et al. (2023); Wu et al. (2022b). Alongside these experimental advancements, theoretical frameworks have been developed for the JDE behavior Cheng and Sun (2023); Davydova et al. (2022a); Hu et al. (2007a); Misaki and Nagaosa (2021b); Lu et al. (2023); Souto et al. (2022); Halterman et al. (2022); Kokkeler et al. (2022); Fominov and Mikhailov (2022); Kopasov et al. (2021); Gupta et al. (2023); Sun et al. (2023); Jiang et al. (2022); Tanaka et al. (2022); Maiani et al. (2023); Hu et al. (2023); Misaki and Nagaosa (2021b); Steiner et al. (2023); Souto et al. (2022); Pillet et al. (2023); Hess et al. (2023); Costa et al. (2025); Fukaya et al. (2024); Debnath and Dutta (2025); Roy et al. (2025); Yerin et al. (2025), of particular interest is the JDE realized in topological nanowires hosting MBSs. Previous studies have focused on ideal one-dimensional (1D) models Liu et al. (2024); Legg et al. (2023); Mondal et al. (2025), whereas realistic systems with multiple subbands Stanescu et al. (2011); Lutchyn et al. (2011); Liu et al. (2012) have been largely overlooked. Such systems may exhibit distinct mechanisms that can significantly modify the diode efficiency.
Figure 1: (a) Quasi-1D Josephson junction nanowire on an $s$-wave SC, with MBSs (blue) and ABSs (light-green). (b) The lowest energy spinful subbands with $(i_{y},i_{z})$ = (1,1) and (2,1) are shown. (c) Energy spectrum in the deep topological region showing isolated MZMs (blue), coupled MZMs (green), and bulk states (black). (d) Interplay of $2\pi$- and $4\pi$-periodic Josephson currents. Dashed blue ($I_{4\pi}$) and green ($I_{2\pi}$) curves denote ground-state MBS and ABS contributions, respectively. The discontinuity in the total current $I_{\mathrm{tot}}$ arises from an abrupt parity switch at the crossing point of the $4\pi$-periodic current $I_{4\pi}$.
@@
The multiband model.–We start with the multiband model for a single quasi-1D nanowire placed on the $s$-wave SC utilizing the proximity effect, forming a Josephson junction as depicted in Fig. 1(a). The system has finite dimensions $L_{x}$, $L_{y}$, $L_{z}$ along the three spatial directions with trapping potentials. In typical nanowire material, the system is strongly confined along the $y$ and $z$ directions, and we consider the rectangular cross section with a width of $L_{y}\sim 10^{2}\,\text{nm}$ and a thickness of $L_{z}\sim 10^{0}\,\text{nm}$, resulting in a reasonable approximation of $L_{z}\ll L_{y}\ll L_{x}$. Consequently, we limit our analysis to the lowest eigenstates in the $z$-direction, setting $N_{z}=1$ Stanescu et al. (2011); Lutchyn et al. (2011). As for the $y$-direction, we focus on cases that a few transverse eigenmodes in this direction can be occupied (e.g. with the number of orbits $N_{y}\leq 3$, see Fig. 1(b)). The resulting effective lattice Hamiltonian takes the form $H_{\mathrm{nw}}=H_{0}+H_{\text{SOI}}+H_{B}+H_{\Delta}$, whose explicit form is given by
@@
with electron creation operator denoted by the spinor $c_{\boldsymbol{i}}^{\dagger}=(c_{\boldsymbol{i}\uparrow}^{\dagger},c_{\boldsymbol{i}\downarrow}^{\dagger})$ and $\boldsymbol{i}=(i_{x},i_{y})$. Note that the index $i_{x}$ denotes the real site position along the $x$-direction, while $i_{y}=1,2,...,N_{y}$ is an effective position. The Hamiltonian consists of four contributions: The free-particle term $H_{0}$ describes nearest-neighbor hopping $t_{x}$ along $\boldsymbol{\delta}_{x}=(1,0)$ and includes an $i_{y}$-site-dependent (actually transverse-mode dependent) chemical potential $\mu_{\boldsymbol{i}}=\mu_{0}-i_{y}^{2}(\pi\hbar)^{2}/(2m^{*}L_{y}^{2})$ Further details are provided in the Supplementary Material , where the energy of transverse modes is approximated by that in a 1D infinite square well Lutchyn et al. (2011), and $m^{*}$ is the effective mass of the electron. The SOI term $H_{\text{SOI}}$ incorporates the longtitutional SOI with intensity $\alpha_{x}$ and the effective transverse SOI between different orbits with intensity $\alpha_{y}\propto L_{y}^{-1}$, which is nonzero only when $\boldsymbol{\delta}_{y}$=(0,1) Further details are provided in the Supplementary Material , with the Pauli matrices $\hat{\sigma}_{i}(i=x,y,z)$ on spin space. The term $H_{B}$ describes a uniform magnetic field $\boldsymbol{h}=(h_{x},h_{y})$ that applied throughout the entire nanowire, where $h_{x}$ opens a Zeeman gap in the spectrum, while $h_{y}$ shifts the Fermi surface along the $x$ direction, breaking inversion symmetry Liu and Lobos (2013); Yuan and Fu (2022a). Finally, $s$-wave pairing is depicted by $H_{\Delta}$, with $\Delta_{\boldsymbol{i}}$ nonzero only in the superconducting leads, carrying a phase difference $\phi$ across the junction: $\Delta_{\mathrm{R}}=\Delta_{\mathrm{L}}e^{i\phi}=\Delta e^{i\phi}$.
Balance between fractional and conventional Josephson currents.– In the multiband regime, the Fermi level intersects several subbands. When the number of Fermi points is odd, the system is topological with MBSs, while the conventional ABSs can coexist at the nanowire ends [Fig. 1(a)]. In a Josephson junction, MBSs couple and contribute a fractional Josephson effect with a $4\pi$ periodicity Kwon et al. (2004b); Fu and Kane (2009); Kwon et al. (2004a); Laroche et al. (2019); Wiedenmann et al. (2016), whereas the conventional ABSs lead to $2\pi$-periodic Josephson currents Sauls (2018); Mizushima and Machida (2018); Hays et al. (2018); Tosi et al. (2019); Nichele et al. (2020). These two components, denoted $I_{4\pi}$ and $I_{2\pi}$ [Fig. 1(d)], respectively, coexist and together determine the total supercurrent at zero temperature as a function of the junction phase difference $\phi$. The total current reads Beenakker (1991)
@@
We adopt experimentally realistic parameters for InAs and InSb nanowires proximitized by Nb or Al superconductors Lutchyn et al. (2011): with $L_{y}=130\,\text{nm}$, $m^{*}=0.04m_{e}$, $\alpha_{x}=0.1\,\text{eV}\cdot\mathring$$\text${A}} and $\Delta=2.4\,\text{K}$ ($0.2\,\text{meV}$). Taking the effective lattice spacing along the $x$ direction to be $a_{x}=4\,\mathrm{nm}$, the corresponding tight-binding parameters are $t_{x}=200\Delta$, $\alpha_{x}=6\Delta$, $\alpha_{y}=1.3\Delta$, and let $\Delta=1$ for convenience. The chemical potential $\mu_{0}$ is set such that the Fermi level crosses an odd number of subbands, ensuring the emergence of MBSs. Moreover, the Fermi level is set to lie midway between the two eigenmodes at $k_{x}=0$. We consider the short junction with a normal region length of $N_{\text{N}}=2$. The Josephson current is evaluated using the recursive Green’s function method Song et al. (2016); Mao et al. (2024b), which efficiently avoids the computational cost in exact diagonalization for large system sizes
@@
Fig. 2 shows the diode efficiency versus $h_{x}$ for different system parameters like the SOI strength $\alpha_{x}$ (Fig. 2(a)), the nanowire’s transverse length $L_{y}$ (Fig. 2(b)), and $h_{y}$ (Fig. 2(c)). For the single band regime (see the colored dashed lines), we can find that the efficiency $\eta$ consistently exhibits a pronounced peak near the critical topological phase transition boundary $h_{x}^{c}=\sqrt{\Delta^{2}-h_{y}^{2}}$ Liu et al. (2024) and diminishes if tuning into deep topological and trivial regimes. In the trivial (topological) regime, the JDE are primarily contributed from the higher harmonic terms of conventional (fractional) Josephson currents $I_{2\pi}$ ($I_{4\pi}$) due to the ABSs (MBSs) Legg et al. (2023), for which the diode effect is weak. Only near the phase transition boundary $h_{x}\approx h_{x}^{c}$, which enables a competition between the two types of Josephson currents, an enhanced JDE is resulted.
@@
Figure 3: High diode efficiency plateau by spin-parity subband exchanges mechanism. (a) Energy bands before, during, and after the spin-parity band exchange in the 2-orbit ($N_{y}=2$) model with 3 occupied subbands. Colored bands denote different spin-parity subbands, and colored arrows mark the Fermi point shifts induced by the inversion-symmetry-breaking field $h_{y}$. (b) Diode efficiency $\eta$ as a function of $h_{x}$, where the background colors correspond to the band structures shown in (a). The current-phase relation at the marked point ($h_{x}=4.8\Delta$) is shown in the insets. (c) Energy bands showing the spin-parity band exchange in the three-orbital ($N_{y}=3$) model with five occupied subbands. (d) $\eta$ versus $h_{x}$, with color segments matching the band structures in (c). Parameters: $\alpha_{x}=10,\,h_{y}=0.8$.
The multiband regime shows a novel new feature that the strong JDE is generically obtained in the deep topological phase regime. Within the multiband topological phase, the contributions from ABSs and MBSs coexist, leading to a persistent competition between $I_{4\pi}$ and $I_{2\pi}$ across the entire topological regime. Consequently, although the maximum position of diode efficiency varies depending on system parameters, the regime of high diode efficiency extends over most of the topological phase region [Fig. 2(a)-(c)].
In Fig. 2(d), increasing the chemical potential drives the Fermi level across one, two, and three subbands, defining regimes I, II, and III. Region $\mathrm{I}$ features a low efficiency platform contributed from MBSs and region $\mathrm{II}$ exhibits weak JDE dominated by ABSs. In contrast, in region $\mathrm{III}$ a high diode efficiency is driven by the coexistence of MBSs and ABSs.
@@
Spin-parity band exchange mechanism.–A more intriguing discovery is that we predict a robust high efficiency plateau with a novel spin-parity band exchange mechanism obtained only in the multiband regime. The spin-parity $\mathbb{P}_{s}$ of a certain subband is defined as the sign of the eigenvalue of the spin-coupled term of the Hamiltonian $h_{x}\hat{\sigma}_{x}+(h_{y}+2\alpha_{x}\sin k_{x}a_{x})\hat{\sigma}_{y}$ at $k_{y}=0$, where the interband SOI vanishes. As illustrated in Fig. 3, the bands with $\mathbb{P}_{s}=+1$ and $\mathbb{P}_{s}=-1$ are plotted in red and blue color, respectively. The relative positions of normal subbands (with $\Delta=0$) with different spin parities can be tuned by $h_{x}$ and shows a crucial impact on the diode effect. As $h_{x}$ increases, the subbands with opposite spin parities shift oppositely in energy, leading to the spin parity exchange of the two types of subbands when $h_{x}$ exceeds a threshold. We consider first the topological phase for $N_{y}=2$, with three normal subbands crossing Fermi energy [Fig. 3(a)]. We can see that at $h_{x}=0.7\Delta$ with $h_{y}=0.8\Delta$, on Fermi level there are two blue bands with $\mathbb{P}_{s}=-1$ and one red band with $\mathbb{P}_{s}=+1$. At $h_{x}\approx|\mu_{1}-\mu_{2}|/2\approx 3.5\Delta$, the upper two subbands begin to swap positions, and the full exchange is completed at $h_{x}\geq 4.7\Delta$. Then the spin parity configuration is robust for further increasing $h_{x}$. The diode efficiency in Fig. 3(b) highlights three distinct regions corresponding to the band exchange shown on the left. Prior to the band exchange, the diode efficiency remains to be low. At the band exchange region, a sharp enhancement in the diode efficiency emerges. After the complete exchange of the two subbands with different spin parities, a robust and high JDE plateau is obtained.
@@
The predicted large and robust JDE is not limited to $N_{y}=2$, but a generic result applicable to the multiband regime. To demonstrate this, we further consider $N_{y}=3$, where the topological phase is realized by tuning the Fermi level to intersect five subbands, comprising two red and three blue bands [Fig. 3(c)]. With increasing $h_{x}$, two spin parity exchange transitions are obtained, respectively at $h_{x}\approx 6.2\Delta$ and $h_{x}\approx 8.8\Delta$. Following the second transition, once the spin-parity configuration becomes stabilized, a robust efficiency plateau generically develops [Fig. 3(d)].
@@
The giant efficiency plateau implies a novel mechanism featured by the spin parity configuration of the subbands. As illustrated in Fig. 3(a), the Fermi level intersects three subbands. For $h_{y}=0$, the bands remain symmetric about $k_{x}=0$. With $h_{y}\neq 0$, the 1D dispersions take the form $E_{i_{y}}^{\pm}(k_{x})=2t_{x}(1-\cos k_{x}a_{x})-\mu_{i_{y}}\pm\sqrt{h_{x}^{2}+(h_{y}+2\alpha_{x}\sin k_{x}a_{x})^{2}},$ with the orbit and spin-parity being labeled by $i_{y}$ and $\pm$, respectively. Therefore, upon applying $h_{y}$, the net shift of the two Fermi points in a given subband is
@@
This expression quantifies the directional deformation of Fermi points, where the red bands ($\mathbb{P}_{s}=+1$) shift leftward and the blue bands ($\mathbb{P}_{s}=-1$) shift rightward [Fig. 3(a)]. Fig. 4(a,b) shows the Fermi point positions and their shifts for the three occupied bands with $N_{y}=2$ as a function of the magnetic field $h_{x}$. We find that $(\Delta k_{\textrm{F}})_{1}^{+}$ and $(\Delta k_{\textrm{F}})_{2}^{-}$ initially decrease, cross zero, and eventually saturate. The zero crossing signifies an exchange of the Fermi points. These Fermi point shifts are induced by the transverse field $h_{y}$. A key feature is that the dimensionless ratio $\xi=2|(\Delta k_{\textrm{F}})_{1}^{+}|/[(\Delta k_{\textrm{F}})_{1}^{-}+(\Delta k_{\textrm{F}})_{2}^{-}]$, which quantifies the relative shift between left- and right-shifted total Fermi momenta, saturates to a stable plateau after the spin-parity exchange [the yellow line in Fig. 4(b)]. This saturation signals the balance between the two types of Josephson currents and corresponds to the emergence of strong JDE. In this regime, the two blue bands with spin-parity $\mathbb{P}_{s}=-1$ exhibit nearly identical and constant Fermi point displacements, which mainly contribute the conventional Josephson current $I_{2\pi}$. In contrast, the remaining red subband predominantly contributes to the fractional Josephson current $I_{4\pi}$ originated from the MBSs.